what is the centripetal acceleration of a point object on the perimeter of a bicycle with radius 35 cm , when the bike is moving a 8 m/s
1
Expert's answer
2020-11-16T07:52:06-0500
"a=\frac{v^2}{r}\\a=\frac{8^2}{0.35}=183\frac{m}{s^2}"